An object moving with constant acceleration has velocity v = 30.8 m/s when it is at x = 3.6 m and v = 55.1 m/s when it is at x = 6.4 m. What is its acceleration? 


1
Expert's answer
2021-02-15T17:44:41-0500

"x_0 = 3.6;V_0=30.8"

"x_1=6.4;V_1=55.1"

"a= \frac{V_1-V_0}{t};t= \frac{V_1-V_0}{a}"

"S_x= x_1-x_0=6.4-3.6=2.8"

"S_x=V_0t+\frac{at^2}{2};"

"S_x=V_0*\frac{V_1-V_0}{a}+\frac{a\frac{(V_1-V_0)^2}{a^2}}{2};"

"S_x=\frac{V^2_1-V^2_0}{2a}"

"a=\frac{V^2_1-V^2_0}{2S_x}=\frac{55.1^2-30.8^2}{2*2.8}=372.2"

Answer: 372.2 m/s2